OpenStudy (australopithecus):

When determining if this integral is convergent or divergent using the f(x) > or = g(x) for all x > or = a \[\int\limits_{0}^{\infty} {xdx}/{(1+x^{3})}\] How do figure out what sign to use? In respect to greater than or equal to or less than or equal to

OpenStudy (australopithecus):

I get that you can split the integral to get a separate integral to do the test \[\int\limits_{0}^{1} xdx/(1+x^{3}) + \int\limits_{1}^{\infty}xdx/(1+x^{3})\] I know the answer is less than or equal to but I dont understand why

OpenStudy (anonymous):

No time to solve this integral, I get http://www.wolframalpha.com/input/?i=integral+xdx%2F%281%2Bx^3%29 And use this tutorial and examples to clarify your doubts http://tutorial.math.lamar.edu/Classes/CalcII/ImproperIntegrals.aspx

OpenStudy (australopithecus):

you dont have to solve this integral you just have to show that it is convergent or divergent

OpenStudy (australopithecus):

using the rule I outlined you create a new function and compare it

OpenStudy (anonymous):

I think I have given the link for tutorial. Have a look at the examples and try

OpenStudy (australopithecus):

Im sorry I dont these problems are related to my question at all I will show you how my prof did it

OpenStudy (australopithecus):

\[x/(1+x^{3}) \le x/(x^{3} + 0)\] since \[\int\limits_{1}^{\infty} (1/x^{2}) dx\] is convergent the original integral is convergent

OpenStudy (anonymous):

@lgbasallote , can you help here please :)

OpenStudy (lgbasallote):

mm sorry i dont know convergent and divergent :C

OpenStudy (anonymous):

Same here. Thanks for coming on my request :)

OpenStudy (lgbasallote):

athough..why is a calculus question in a physics group :p haha

OpenStudy (lgbasallote):

im sure @amistre64 can help..he's good with calculus :)

OpenStudy (anonymous):

@amistre64 ,calculus help bro. Hope you can do it :D

OpenStudy (amistre64):

are we asking why a contiuous sum is greater than a discrete sum?

OpenStudy (amistre64):

imagine it like this; line up a bunch of pennies so they form a contiuous line; count them. then remove every other penny so that there is a gap between them; then add the up again the continuous line is greater in value that the discrete line

OpenStudy (amistre64):

if the sum of the greater converges, the lesser has to converge

OpenStudy (australopithecus):

still confused but I will try to comprehend what you are saying :L